OpenStudy (dmezzullo):

Compare and Contrast. Below are two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set #1 Set #2 (1, 5) and (2, 9) (6, 2) and (3, 4) The slope of Set #1 is smaller than the slope of Set #2. The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is the same as the slope of Set #2. None of the statements above describes slopes of sets shown.

OpenStudy (anonymous):

What is the formula for finding slope?

OpenStudy (dmezzullo):

errmm idk

OpenStudy (anonymous):

You cant do math without formulas. Check your textbook. Its a pretty easy question, if you know the formula.

OpenStudy (anonymous):

y= mx+b

OpenStudy (dmezzullo):

Yes wut she said

OpenStudy (anonymous):

y=mx+b has got no use here. Slope=(y2-y1)/(x2-x1)

OpenStudy (dmezzullo):

ok

OpenStudy (anonymous):

i didnt know which one it was but yeah rnr's right

OpenStudy (dmezzullo):

wait wut?

OpenStudy (anonymous):

Find the slopes and compare. You just need to plug in the values now.

OpenStudy (dmezzullo):

Would it be B?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

True.

OpenStudy (anonymous):

awh!!!! ok and thanks fer the medal!!

OpenStudy (dmezzullo):

Np Madam